what is the area of a right triangle with an 45 and one leg has a length of 13
Answer by nerdybill(7384)

what is the area of a right triangle with an 45 and one leg has a length of 13
.
Since, it is 45, we know it is an isosceles triangle.
therefore,
area = (1/2)bh
area = (1/2)(13)(13)
area = (1/2)(169)
area = 84.5